in case of projectile fired at angle equally inclined to th e horizontal and vertical with velocity u , the horizotal range is given by
 

Grade:8

1 Answers

Mahima Kanawat
6 years ago
Here , i am giving the sign © to theta R= u^2sin2©/g Here © = 45° [equally inclined ]ie., sin2©=sin2(45°)=sin90°=1ie., R = u^2/g